# A unified weighting framework for evaluating nearest neighbour classification

### v2

@@ 1 Introduction
Nearest neighbours (NN) [1] is one of the oldest and most widely used classification algorithms. Despite the development of other successful classifiers, it remains an attractive option in many settings due to its simplicity [2] and good performance for many problems, especially when decision boundaries are irregular [3]. NN and its variants are applied in such diverse domains as healthcare [4, 5, 6], agriculture [7, 8], geology [9, 10], industry [11, 12, 13] and natural language processing [14, 15, 16], and have also been adapted to big data settings [17, 18, 19].
While NN is a relatively simple algorithm, it still requires setting a few hyperparameters. Even in its most basic form, we need to choose a distance measure and a method to rescale the attributes of the dataset. In addition, it is generally advisable to choose the number of neighbours $k\geq 1$ on which predictions are based. Moreover, there have been many proposals in the literature to weigh the contribution of the $k$ nearest neighbours of a test record differently.
@@ 2.1 Weighted nearest neighbour classification
Finally, Gou et al. [32] have proposed a weighting scheme that combines the linear distance-weights of Dudani with reciprocal rank-weights:
@@ 2.3 Fuzzy nearest neighbour classification
There have been many proposals to modify nearest neighbour classification with fuzzy set theory [42]. The most prominent of these is the fuzzy nearest neighbours (FNN) classifier (Fig. 2) of Keller et al. [20]. It defines the membership of a test record $y$ in the decision class $C$ as
@@ 2.4 Fuzzy rough nearest neighbour classification
A more fundamentally different proposal has come in the form of fuzzy rough nearest neighbour (FRNN) classification (Fig. 3), originally proposed by Jensen & Cornelis [21]. This is based on fuzzy rough sets [43], a fuzzified variant of rough sets [44]. For each decision class $C$, we define two fuzzy sets, its upper approximation $\overline{C}$ and its lower approximation $\underline{C}$, as well as their mean, and the membership of a test record $y$ in any one of these can be used as a class score. A weighted variant of fuzzy rough sets was first introduced by Cornelis et al. [45], and we use here the updated formulation of FRNN presented in [18].
@@ 3.1 Kernels
In the context of nearest neighbour weighting, all ranks and distance values are non-negative, and the weights are distributed over a finite range, limited by the $k$th rank or nearest neighbour distance. Therefore, all weight types that we will consider can be covered by the following more restricted definition of a kernel:
@@
A (proper) kernel is a decreasing function $f:[0,1]\longrightarrow\mathbb{R}_{\geq 0}$. A normalised kernel is a kernel with $f(0)=1$. An improper kernel is a decreasing function $f:(0,1]\longrightarrow\mathbb{R}_{\geq 0}$ with $\lim_{a\rightarrow 0}f(a)=\infty$. For a proper or improper kernel $f$, we write $f^{k}_{i}=f(\frac{i}{k+1})$.
@@
In the following subsections, we will redefine NN and FRNN classification in terms of kernel functions. The kernels used in this paper are listed in Table 1 and the proper kernels are visualised in Fig. 4.
@@
$\left.\sum_{i\leq k|x_{i}\in C}w_{i}^{k}\cdot s(d_{i}^{*})\middle/\left(\sum_{i\leq k}w_{i}^{k}\cdot s(d_{i}^{*})\right)\right.,$
where $x_{i}$ is the $i$th nearest neighbour of $y$ in the training set $X$ (as determined by $d$), $d_{i}$ the corresponding distance and $d_{i}^{*}:=d_{i}/d_{k}$.
@@
We will now show how this incorporates existing weighting proposals. Firstly, note that both in the original equation for NN (1), and in (3), we rescale each class score by the total sum of the weights. Therefore, we do not require that weights sum to 1. Moreover, multiplying all weights by a positive constant produces identical class scores. We will use the proportionality symbol $\propto$ to indicate that two sets of weights only differ by a positive constant factor. Thus, the linear distance-weights proposed by Dudani [28] can be simplified as follows:
$s_{i}$ $=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}$
$\propto\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}-d_{1}}{d_{k}}$
$=1-d_{i}^{*}.$
We see that Dudani’s linear weights are equivalent to applying a linear (or triangular) kernel in our revised definition. In fuzzy set theory, this is the original fuzzy negation introduced by both Łukasiewicz [54] and Zadeh [55].
In a similar way, we can simplify the weight types proposed by Gou et al. [23]:
$s_{i}$ $=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}+d_{1}}{d_{k}+d_{i}}$ (6)
$=\frac{1-d_{i}^{*}}{1+d_{i}^{*}}.$
We will call this the Sugeno-1 kernel, because it is in fact an instance of fuzzy Sugeno negation [56], with $\lambda=1$:
$a\longmapsto\frac{1-a}{1+\lambda a}$
Weights that depend reciprocally linearly or reciprocally quadratically on nearest neighbour distance result from a straightforward application of an improper kernel, since, respectively,
$s_{i}=\frac{1}{d_{i}}\propto\frac{d_{k}}{d_{i}}=\frac{1}{d_{i}^{*}},$
and
$s_{i}=\frac{1}{d_{i}^{2}}\propto\frac{d_{k}^{2}}{d_{i}^{2}}=\frac{1}{\left(d_{i}^{*}\right)^{2}}.$
We can also simplify the weight types proposed by Macleod et al. [37]:
$s_{i}$ $=\frac{d_{k}-d_{i}+d_{k}-d_{1}}{2(d_{k}-d_{1})}$
$=2-d_{i}^{*}-d_{1}^{*}.$
Similarly, most types of rank-weights proposed in the literature can be obtained by applying a kernel function to $\frac{i}{k+1}$. For the Samworth weights, if we fix a particular value $k>0$, and write $h=\frac{1}{k+1}$ (for reasons of space), we can define the following kernel function $f^{k}$ such that $f^{k}(\frac{i}{k+1})$ is the $i$th weight as in (2):
Lemma 1.
$\lim_{k\rightarrow\infty}\frac{2k}{m+2}\cdot f^{k}=1-a^{\frac{2}{m}}.$
Proof.
$\mathrel{\phantom{=}}\lim_{k\rightarrow\infty}\frac{2k}{m+2}\cdot f_{k}$
$=1-\frac{m}{m+2}\cdot\left(1+\frac{2}{m}\right)a^{\frac{2}{m}}$ (*)
$=1-a^{\frac{2}{m}},$
where ()(*)( ∗ ) is the polynomial rule for derivation. ∎
Accordingly, we call $1-a^{\frac{2}{m}}$ the Samworth kernel.
The only rank-weights that can not be obtained by applying a kernel function are the Fibonacci weights from [33], because their relative distribution depends on $k$.
@@ 3.3 Yager weights
In the previous section we found that two weighting proposals from the literature correspond, respectively, to classical and Sugeno negation. There is a third type of fuzzy negation that is frequently encountered in the literature, with the following form:
@@ 3.4 FNN
3.5 FRNN
Definition 4.
[…]
where $d^{+}_{i}$ and $d^{-}_{i}$ are the $i$th nearest neighbour distance of $y$ in, respectively, $C$ and $X\setminus C$, and $d^{+}_{*}$ and $d^{-}_{*}$ are to be defined.
3.6 Distance and scaling measures
Definition 5.
@@ 4 Experimental setup
To evaluate NN, FNN and FRNN classification, we will use 85 numerical real-life datasets from the UCI repository for machine learning (Table 2). We perform 5-fold cross-validation, and calculate the mean area under the receiver operating characteristic (AUROC) as a measure of the discriminative ability of each classifier. For FRNN, this requires normalising class scores such that they sum to 1. To compare two alternatives, we calculate the $p$-value from a one-sided Wilcoxon signed-rank test. Where appropriate, we also apply the Holm-Bonferroni method [58] to correct for family-wise error.
@@ 5 Results
5.1 NN
Table 3: One-sided $p$-values, Samworth distance-weights vs other distance-weights, for NN with Boscovich distance, in terms of AUROC. Holm-Bonferroni family-wise error correction applied to each column.
@@
Biquadratic $0.014$ $0.11$ $0.71$ $0.042$
Reciprocally linear $0.0012$ $0.00016$ $<0.0001$ $<0.0001$
Reciprocally quadratic $0.038$ $0.020$ $0.046$ $0.059$
Sugeno-1 $0.0068$ $0.0036$ $0.029$ $0.00036$
We first consider the effect of the distance on classification performance. We find that for all types of scaling and all weight types, Boscovich distance leads to significantly better performance than Euclidean ($p<0.0031$) and Chebyshev ($p<2.0\cdot 10^{-9}$) distance. For this reason, we will only consider Boscovich distance for the rest of our analysis.
Next, we have a look at the different kernels. First we compare using each kernel for distance-weights versus rank-weights, for each type of scaling. With three exceptions, we find that distance-weights lead to significantly better classification performance than rank-weights ($p<0.0019$). The exceptions are the reciprocal and Laplace kernels, for which the difference is not or only weakly significant and to which we return below, as well as the Gaussian kernel, for which the difference is only weakly significant with $r_{\infty}$-scaling ($p=0.26$).
Among the distance-weights, the Samworth kernel significantly outperforms all other kernels (Table 3), with three exceptions. The difference with respect to the biquadratic and reciprocally quadratic kernel is only weakly significant with, respectively, $r_{2}$- and $r_{\infty}^{*}$-scaling. Moreover, with $r_{\infty}$-scaling, the biquadratic kernel is actually slightly better than the Samworth kernel on our data, but we will see below that $r_{\infty}$-scaling is suboptimal. Finally, we also find that Samworth distance-weights outperform reciprocal ($p<0.0048$) and Laplace ($p<6.2\cdot 10^{-6}$) rank-weights across scaling measures.
@@
Table 4: One-sided $p$-values, various scalings vs $r_{\infty}$-scaling, for NN with Boscovich distance and Samworth distance- and rank-weights, in terms of AUROC. Holm-Bonferroni family-wise error correction applied.
@@ 5.2 FNN
Unlike NN, it is not clear that Samworth distance-weights perform better than reciprocally linear or reciprocally quadratic weights (Table 5). Furthermore, the combination of Samworth rank- and distance-weights actually performs worse than Samworth distance-weights alone for $r_{1}$ ($p=0.032$) and $r_{\infty}$ ($p=0.000078$) scaling, while for $r_{2}$- and $r_{\infty}^{*}$-scaling, the difference is not significant. We have weak evidence that with FNN, $r_{1}$-scaling leads to better performance than $r_{2}$-scaling, and that in turn both are preferable over $r_{\infty}$- and $r_{\infty}^{*}$-scaling (Table 6).
Table 5: One-sided $p$-values, Samworth distance-weights vs other distance-weights, for FNN with Boscovich distance, in terms of AUROC. Holm-Bonferroni family-wise error correction applied to each column.
@@
Table 6: One-sided $p$-values, comparing the scaler in each row against the scaler in each column, for FNN with Boscovich distance and Samworth distance-weights, in terms of AUROC. Holm-Bonferroni family-wise error correction applied to each row.
@@
Constant $0.18$ $0.0033$ $0.089$ $0.0048$
Linear $0.18$ $0.074$ $0.73$ $0.32$
Reciprocally linear $0.18$ $0.022$ $0.15$ $0.057$
@@ 5.3 FRNN
Table 9: One-sided $p$-values, various scalings vs $r_{\infty}$-scaling, for FRNN with Boscovich distance and Samworth distance- and rank-weights, in terms of AUROC. Holm-Bonferroni family-wise error correction applied.
@@
$r_{1}$ vs $r_{\infty}$ $0.060$
$r_{\infty}^{*}$ vs $r_{\infty}$ $0.060$
For FRNN, we evaluate the different types of rank-weights proposed in the literature, corresponding to the constant, linear and reciprocal kernel, in combination with linear distance-weights. In addition, we evaluate Samworth rank- and distance-weights, motivated by their excellent performance with NN.
Here we also find that Boscovich distance leads to higher AUROC than Euclidean ($p<0.00085$) and Chebyshev ($p<6.2\cdot 10^{-10}$) distance for all combinations of kernels and scaling measures.
The traditional choice for distance-weights is to use a linear kernel, but we find that the Samworth kernel performs better, although the difference is only weakly significant in combination with reciprocal rank-weights (Table 7).
Likewise, Samworth rank-weights appear to be the best choice, but the advantage over other kernels is only weakly significant (Table 8).

### v3

@@ 1 Introduction
Nearest neighbours (NN) [1] is one of the oldest and most widely used classification algorithms. It identifies the nearest neighbours (the most similar records) of a given test record in the training set, and predicts class scores based on the prevalence of each decision class among these neighbours. Despite the development of other successful classifiers, it remains an attractive option in many settings due to its simplicity [2] and good performance for many problems, especially when decision boundaries are irregular [3]. Even when other algorithms achieve higher classification performance, NN may be preferable due to its interpretability: every prediction can be explained to a practitioner by comparing a test record with a handful of its nearest neighbours. NN and its variants are applied in such diverse domains as healthcare [4, 5, 6], agriculture [7, 8], geology [9, 10], industry [11, 12, 13] and natural language processing [14, 15, 16], and have also been adapted to big data settings [17, 18, 19].
While NN is a relatively simple algorithm, it still requires setting a few hyperparameters. Even in its most basic form, we need to choose a distance measure and a method to rescale the attributes of the dataset. In addition, it is generally advisable to choose the number of neighbours $k\geq 1$ on which predictions are based. Choosing a larger value for $k$ increases the bias of the model and its tolerance for noisy training records, at the cost of reduced flexibility. Finally, there have been many proposals in the literature to weigh the contribution of the $k$ nearest neighbours of a test record differently.
@@ 2.1 Weighted nearest neighbour classification
Finally, Gou et al. [32] have proposed a weighting scheme that combines the linear distance-weights of Dudani with reciprocally linear rank-weights:
@@ 2.3 Fuzzy nearest neighbour classification
There have been many proposals to modify nearest neighbour classification with fuzzy set theory [42]. The most prominent of these is the fuzzy nearest neighbours (FNN) classifier (Fig. 2) of Keller et al. [20]. It defines the membership of a test record $y$ in the decision class $C$ as:
@@ 2.4 Fuzzy rough nearest neighbour classification
A more fundamentally different proposal has come in the form of fuzzy rough nearest neighbour (FRNN) classification (Fig. 3), originally proposed by Jensen & Cornelis [21]. This is based on fuzzy rough sets [43], a fuzzified variant of rough sets [44]. For each decision class $C$, we define two fuzzy sets, its upper approximation $\overline{C}$ and its lower approximation $\underline{C}$, as well as their mean, and the membership of a test record $y$ in any one of these can be used as a class score. A weighted variant of fuzzy rough sets was first introduced by Cornelis et al. [45], and we use here the updated formulation of FRNN presented in [18]. FRNN can be viewed as a form of nearest neighbour classification because the membership of $y$ in $\overline{C}$ and $\underline{C}$ is defined as a weighted sum of, respectively, its nearest neighbour similarities in $C$, and its nearest neighbour distances in $X\setminus C$.
@@ 3.1 Kernels
In the context of nearest neighbour weighting, all ranks and distance values are non-negative, and the weights are distributed over a finite range, limited by the $k$th rank or nearest neighbour distance, which can be rescaled to the interval $[0,1]$. Therefore, all weight types that we will consider can be covered by the following more restricted definition of a kernel:
@@
A (proper) kernel is a decreasing function $f:[0,1]\longrightarrow\mathbb{R}_{\geq 0}$. A normalised kernel is a kernel with $f(0)=1$. An improper kernel is a decreasing function $f:(0,1]\longrightarrow\mathbb{R}_{\geq 0}$ with $\lim_{a\rightarrow 0}f(a)=\infty$.
@@
In the following subsections, we will redefine NN, FNN and FRNN classification in terms of kernel functions, and show how the various weighting proposals from the literature can be expressed through kernels by rescaling the distance and rank values to the interval $[0,1]$. The advantage of this approach is that, for most weight types, this eliminates the dependence of the weights on $k$, i.e. they can be expressed as a single kernel function that is ‘stretched’ across the rank or distance values. [^1] For linear weights, this is easy to see as the slope of the weights depends on $k$. Thus, the kernel function captures the essence of a weighting type, and comparing these kernel functions reveals the essential differences between the various weighting proposals. The kernels discussed in this paper are listed in Table 1 and the proper kernels are visualised in Fig. 4.
@@
$\left.\sum_{i\leq k|x_{i}\in C}w(i^{*})\cdot s(d_{i}^{*})\middle/\left(\sum_{i\leq k}w(i^{*})\cdot s(d_{i}^{*})\right)\right.,$ (6)
where $x_{i}$ is the $i$th nearest neighbour of $y$ in the training set $X$ (as determined by $d$), $d_{i}$ the corresponding distance, $d_{i}^{*}:=d_{i}/d_{k}$ and $i^{*}:=\frac{i}{k+1}$.
@@
In A, we show that most existing weighting proposals discussed in Subsection 2.1 can be rewritten as kernel functions (Table 1, Fig. 4) applied to $i^{*}$ (rank-weighting) or $d_{i}^{*}$ (distance-weighting), making them compatible with our generalised definition of NN (6). In particular, note that the linear (or triangular) kernel is the original form of fuzzy negation introduced by both Łukasiewicz [54] and Zadeh [55], that the kernel corresponding to the weight type proposed by Gou et al. [23] is Sugeno-$\lambda$ negation with $\lambda=1$, and that the rank-weights shown to be theoretically optimal by Samworth [22] correspond to a kernel $a\mapsto 1-a^{\frac{2}{m}}$, where $m$ is the dimensionality of the attribute space.
@@ 3.3 Yager weights
As mentioned in the previous subsections, kernels are slightly generalised fuzzy negation functions, and two fuzzy negations from the fuzzy logic literature correspond to existing weighting proposals: classical (linear) and Sugeno-$\lambda$ negation. There is in fact a third type of fuzzy negation that is frequently encountered in the literature, with the following form:
@@ 3.4 FNN
3.5 FRNN
Definition 4.
[…]
where $i^{*}:=\frac{i}{k+1}$, where $d^{+}_{i}$ and $d^{-}_{i}$ are the $i$th nearest neighbour distance of $y$ in, respectively, $C$ and $X\setminus C$, and where $d^{+}_{*}$ and $d^{-}_{*}$ are to be defined.
3.6 Distance and scaling measures
The parameter $p$ determines the contribution of the per-attribute differences $y_{i}-x_{i}$ to the distance value. When $p=1$ (Boscovich distance), the contribution of all per-attribute differences is equal, whereas when $p\rightarrow\infty$ (Chebyshev distance), the distance is simply equal to the largest per-attribute difference, and all smaller per-attribute differences make no contribution. This means that with Chebyshev distance, two records are only neighbours if all their attribute values are similar, whereas with Boscovich distance, it is sufficient that their attribute values are similar on average.
The case $p=2$ (Euclidean distance) stakes a middle ground between these two extremes, in that smaller per-attribute differences make less than a proportional contribution to the distance, without being completely ignored. A unique property of Euclidean distance is that it is invariant under rotations of the attribute space. Therefore, Euclidean distance is the appropriate distance measure when applying a rotation would result in an equivalent representation of the data.
Definition 5.
@@ 4 Experimental setup
To evaluate NN, FNN and FRNN classification, we will use 85 numerical real-life datasets from the UCI repository for machine learning (Table 2). We perform 5-fold cross-validation, and calculate the mean area under the receiver operating characteristic (AUROC) as a measure of the discriminative ability of each classifier. For FRNN, this requires normalising class scores such that they sum to 1. To compare two alternatives, we calculate the $p$-value from a one-sided Wilcoxon signed-rank test. Where appropriate, we also apply the Benjamini-Hochberg method [57] to control the false discovery rate.
@@ 5 Results
Table 3: One-sided $p$-values, Samworth distance-weights vs other distance-weights, for NN with Boscovich distance, in terms of AUROC. With Benjamini-Hochberg false discovery rate correction in each column.
@@
Biquadratic $0.0075$ $0.11$ $0.71$ $0.024$
Reciprocally linear $0.00044$ $<0.0001$ $<0.0001$ $<0.0001$
Reciprocally quadratic $0.038$ $0.011$ $0.025$ $0.059$
Sugeno-1 $0.0029$ $0.0015$ $0.012$ $0.00015$
5.1 NN
We first consider the effect of the distance on classification performance. We find that for all types of scaling and all weight types, Boscovich distance leads to significantly better performance than Euclidean ($p<0.0031$) and Chebyshev ($p<2.0\cdot 10^{-9}$) distance. For this reason, we will continue with Boscovich distance only. However, we investigate this question in greater detail in Subsection 5.5.
Next, we have a look at the different kernels. First we compare using each kernel for distance-weights versus rank-weights, for each type of scaling. With three exceptions, we find that distance-weights lead to significantly better classification performance than rank-weights ($p<0.0019$). The exceptions are the reciprocally linear and Laplace kernels, for which the difference is not or only weakly significant and to which we return below, as well as the Gaussian kernel, for which the difference is only weakly significant with $r_{\infty}$-scaling ($p=0.26$).
Among the distance-weights, the Samworth kernel significantly outperforms all other kernels (Table 3), with three exceptions. The difference with respect to the biquadratic and reciprocally quadratic kernel is only weakly significant with, respectively, $r_{2}$- and $r_{\infty}^{*}$-scaling. Moreover, with $r_{\infty}$-scaling, the biquadratic kernel is actually slightly better than the Samworth kernel on our data, but we will see below that $r_{\infty}$-scaling is suboptimal. Finally, we also find that Samworth distance-weights outperform reciprocally linear ($p<0.0048$) and Laplace ($p<6.2\cdot 10^{-6}$) rank-weights across scaling measures.
@@
Table 4: One-sided $p$-values, various scalings vs $r_{\infty}$-scaling, for NN with Boscovich distance and Samworth distance- and rank-weights, in terms of AUROC. With Benjamini-Hochberg false discovery rate correction.
@@ 5.2 FNN
Table 5: One-sided $p$-values, Samworth distance-weights vs other distance-weights, for FNN with Boscovich distance, in terms of AUROC. With Benjamini-Hochberg false discovery rate correction in each column.
@@
Table 6: One-sided $p$-values, comparing the scaler in each row against the scaler in each column, for FNN with Boscovich distance and Samworth distance-weights, in terms of AUROC. With Benjamini-Hochberg false discovery rate correction in each row.
@@
Constant $0.080$ $0.0033$ $0.089$ $0.0048$
Linear $0.080$ $0.074$ $0.73$ $0.32$
Reciprocally linear $0.080$ $0.016$ $0.11$ $0.043$
However, it is not as clear as with NN that Samworth distance-weights perform better than reciprocally linear or reciprocally quadratic weights (Table 5). Furthermore, the combination of Samworth rank- and distance-weights actually performs worse than Samworth distance-weights alone for $r_{1}$ ($p=0.032$) and $r_{\infty}$ ($p=0.000078$) scaling, while for $r_{2}$- and $r_{\infty}^{*}$-scaling, the difference is not significant. We have weak evidence that with FNN, $r_{1}$-scaling leads to better performance than $r_{2}$-scaling, and that in turn both are preferable over $r_{\infty}$- and $r_{\infty}^{*}$-scaling (Table 6).
@@ 5.3 FRNN
Table 9: One-sided $p$-values, various scalings vs $r_{\infty}$-scaling, for FRNN with Boscovich distance and Samworth distance- and rank-weights, in terms of AUROC. With Benjamini-Hochberg false discovery rate correction.
@@
For FRNN, we evaluate the different types of rank-weights proposed in the literature, corresponding to the constant, linear and reciprocally linear kernel, in combination with linear distance-weights. In addition, we evaluate Samworth rank- and distance-weights, motivated by their excellent performance with NN.
As with NN and FNN, we find that Boscovich distance leads to higher AUROC than Euclidean ($p<0.00085$) and Chebyshev ($p<6.2\cdot 10^{-10}$) distance for all combinations of kernels and scaling measures.
The traditional choice for distance-weights is to use a linear kernel, but we find that the Samworth kernel performs better, although the difference is only weakly significant in combination with reciprocally linear rank-weights (Table 7). Likewise, Samworth rank-weights appear to be the best choice, but the advantage over other kernels is only weakly significant (Table 8).
@@ 5.4 NN vs FNN vs FRNN
5.5 Distance measures
At the start of Subsections 5.1 and 5.3, we found that, regardless of weights and scaling, NN and FRNN generally perform better with Boscovich distance than with Euclidean and Chebyshev distance. We now consider whether, despite this general trend, there might be types of datasets for which Euclidean and Chebyshev distance present an advantage. [^2] Note that we obtain similar results regarding the performance of different weight types and scaling measures for Euclidean and Chebyshev distance as we did for Boscovich distance.
In Fig.6, we compare the performance of NN and FRNN with Boscovich distance against Euclidean and Chebyshev distance. For datasets above the baseline, classification performance is higher for Boscovich distance, whereas for datasets below the baseline, performance is higher for Euclidean or Chebyshev distance. The x-axis expresses how difficult the classification problem is.
The datasets below the baseline, for which Euclidean or Chebyshev distance outperforms Boscovich distance, mostly fall into two subgroups. The datasets on the left side of the plot are very difficult to classify, which introduces a degree of randomness, making even substantial performance differences between distance measures not very meaningful. The datasets on the right side of the plot below the baseline are still very close to it, meaning that the performance difference is very small. Overall, we are not able to identify a specific dataset type for which Euclidean or Chebyshev distance possesses a specific advantage.
On the other hand, we note that Chebyshev distance in particular does lead to substantially poorer performance on the rejafada, tcga-pancan-hiseq, androgen-receptor and oral-toxicity datasets. These are the four datasets in our collection with the highest dimensionality. Because the Chebyshev distance between two records is completely determined by the attribute with the highest difference, it effectively discards a large part of the information contained in these datasets.
Figure 6: AUROC difference (y-axis) between Boscovich distance and Euclidean (top row) or Chebyshev (bottom row) distance, offset against the maximum of the two AUROC scores (x-axis), for NN (left column) and FRNN (right column), with Samworth distance- and rank-weights and $r_{2}$-scaling, applied to the datasets in this paper.